# CogCoM: Train Large Vision-Language Models Diving into Details through Chain of Manipulations

### v2

@@ Appendix C Details of Data Production
In this section, we further introduce the details of CoM data production, with the overall algorithm of a pseudo code, an example of the solving steps generation with LLM and corresponding guideline, an example of the reasoning chains completion with visual tools. We also list the details of data statistics for the synthesised training data as well as the evaluation data of CoM-test, followed by a limitation analysis for the current data production method.
C.1 Algorithm for the Automate Data Generation Pipeline
@@
C.2 The CoM-test Benchmark and Evaluation Metric
To measure the correctness of CoM chains, we introduce a keypoints-aware metric. The intuition is that we care about the key elements including actions (i.e., manipulation name), targets (i.e., manipulation input), and visual contents (i.e., manipulation returns) of each step in the path, as well as the logical execution order of manipulations. Given a pair of chain-answer annotation $(c,a)$ and corresponding model prediction $(c^{\prime},a^{\prime})$, we first sequentially extract the key elements from $c$ and $c^{\prime}$ to construct two ordered lists, and then replace the elements in the lists with their fixed indices in a Bag-of-Elements $\mathcal{E}=c\cup c^{\prime}$ to result in lists of $k$ and $k^{\prime}$. We thus calculate the score as the normalized Levenshtein Distance $s_{c}=Levenshtein(k,k^{\prime})/N$ between the two lists, where $N$ is the maximum length between $k$ and $k^{\prime}$. We adopt this simple discretization strategy with low time complexity to concentrate on the key points as well as the solving order. We further consider the linguistic matching of paragraphs by calculating the BLEU (Papineni et al., 2002) score between two chains $s_{p}=\text{BLEU}(c,c^{\prime})$, and the final sore is a weighted combination as $acc=(0.6\times s_{c}+0.4\times s_{p})/2$.
C.3 Data Statistics
We develop a strategy to extract predicate phrases based constituency parsing with StandordCoreNLP, in which we extract verb, conjunction-connected verb phrase, preposition-connected verb phrase.
Besides the standard CoM data incorporating manipulations with explicit visual evidences, the proposed data synthesising framework is compatible of producing implicit visual reasoning steps $step^{\prime}_{i}=(desc_{i})$ without involving the manipulations. We thereby also build this partial CoM data on the corpus consisting of absurd visual questions (i.e., asking unanswerable questions based on the given image) to further resist the toxic hallucinations. Specifically, given an image $I$ with a question $Q$,we prompt GPT-4V (OpenAI, 2023b) to solve the question step-by-step to acquire the reasoning chains.
@@ C.3 Data Statistics
C.4 Details of the Linguistic/Visual Annotations
In this work, we adopt the GPT4-turbo as the linguistic annotator for generating problems solving steps, and the API call was conducted during the period of 2023.9 - 2023.12. For the visual annotators, we leverage the the currently best-performing tools, GroundingDINO and PaddleOCR, to acquire all visual contents requested by the manipulations. For a clear description to the production setting and results, we illustrate the guiding prompt, and an example-based linguistic annotation results as well as the visual annotation results in Figure 8.
C.5 Limitation Analysis for the Data Production
@@
Appendix D Details of Training
D.1 Launching Prompts
@@ D.2 Training settings
Appendix E Details of Qualitative Analysis
E.1 Qualitative Analysis
We investigate the evidential reasoning capability of CogCoM on scenarios that requires different types of meticulous reasoning, including recognizing textual details, reading time, understanding charts and counting objects. The results are shown in Figure 1. The first case demonstrates that CogCoM finds the region corresponding to the plane logo through two steps of grounding and then achieves the answer based on zooming in the cropped region. The second case illustrates the ability of CogCoM in reading time, by locating the device that displays time and then transforming the time into words based on the read_timne manipulation. In the forth example, CogCoM first identifies all visible truck wheels through grounding, and then leverages the counting manipulation to consolidate the total number.
E.2 Representative Cases
Figure 9: CogCoM demonstrates the flexible capabilities for adapting to different multimodal scenarios, including evidential visual reasoning, Visual Grounding, Grounded Captioning, Image Captioning, Multi Choice, and Detailed Captioning.

### v3

@@ Appendix E Details of Data Production
In this section, we further introduce the details of CoM data production, with the overall algorithm of a pseudo code, an example of the solving steps generation with LLM and corresponding guidelines, an example of the reasoning chains completion with visual tools. We also list the details of data statistics for the synthesized training data as well as the evaluation data of CoM-test, followed by a limitation analysis for the current data production method.
E.1 Algorithm for the Automate Data Generation Pipeline
@@
E.2 The CoM-test Benchmark and Evaluation Metric
To measure the correctness of CoM chains, we introduce a keypoints-aware metric. The intuition is that we care about the key elements including actions (i.e., manipulation name), targets (i.e., manipulation input), and visual contents (i.e., manipulation returns) of each step in the path, as well as the logical execution order of manipulations. Given a pair of chain-answer annotation $(c,a)$ and corresponding model prediction $(c^{\prime},a^{\prime})$, we first sequentially extract the key elements from $c$ and $c^{\prime}$ to construct two ordered lists, and then replace the elements in the lists with their fixed indices in a Bag-of-Elements $\mathcal{E}=c\cup c^{\prime}$ to result in lists of $k$ and $k^{\prime}$. We thus calculate the score as the normalized Levenshtein Distance $s_{c}=Levenshtein(k,k^{\prime})/N$ between the two lists, where $N$ is the maximum length between $k$ and $k^{\prime}$. We adopt this simple discretization strategy with low time complexity to concentrate on the key points as well as the solving order. We further consider the linguistic matching of paragraphs by calculating the BLEU (Papineni et al., 2002) score between two chains $s_{p}=\text{BLEU}(c,c^{\prime})$, and the final score is a weighted combination as $acc=(0.6\times s_{c}+0.4\times s_{p})/2$.
E.3 Data Statistics
We develop a strategy to extract predicate phrases based constituency parsing with StandordCoreNLP, in which we extract verb, conjunction-connected verb phrase, preposition-connected verb phrases.
Besides the standard CoM data incorporating manipulations with explicit visual evidence, the proposed data synthesizing framework is compatible of producing implicit visual reasoning steps $step^{\prime}_{i}=(desc_{i})$ without involving the manipulations. We thereby also build this partial CoM data on the corpus consisting of absurd visual questions (i.e., asking unanswerable questions based on the given image) to further resist the toxic hallucinations. Specifically, given an image $I$ with a question $Q$,we prompt GPT-4V (OpenAI, 2023b) to solve the question step-by-step to acquire the reasoning chains.
@@ E.3 Data Statistics
E.4 Details of the Linguistic/Visual Annotations
In this work, we adopt the GPT4-turbo as the linguistic annotator for generating problems-solving steps, and the API call was conducted during the period of 2023.9 - 2023.12. For the visual annotators, we leverage the currently best-performing tools, GroundingDINO and PaddleOCR, to acquire all visual contents requested by the manipulations. For a clear description for the production setting and results, we illustrate the guiding prompt, and an example-based linguistic annotation results as well as the visual annotation results in Figure 9.
E.5 Limitation Analysis for the Data Production
@@
Appendix F Details of Training
F.1 Launching Prompts
@@ F.2 Training settings
Appendix G Details of Qualitative Analysis
G.1 Qualitative Analysis
We investigate the capabilities of CogCoM on scenarios that requires different types of detailed reasoning, including recognizing textual details, reading time, understanding charts and counting objects. The results are shown in Figure 1. The first case demonstrates that CogCoM finds the region corresponding to the plane logo through two steps of grounding and then achieves the answer based on zooming in the cropped region. The second case illustrates the ability of CogCoM in reading time, by locating the device that displays time and then transforming the time into words based on the read_timne manipulation. In the fourth example, CogCoM first identifies all visible truck wheels through grounding, and then leverages the counting manipulation to consolidate the total number.
G.2 Representative Cases
Figure 10: CogCoM demonstrates flexible capabilities for adapting to different multimodal scenarios, including evidential visual reasoning, Visual Grounding, Grounded Captioning, Image Captioning, Multi Choice, and Detailed Captioning.